OpenStudy (monkey*):

X-y+z=5 -2x-y-z=-6 3x+3y-2z=-5 Solve algebraically.

OpenStudy (texaschic101):

ohhhh...these are a pain in the butt. Very time consuming. here we go... x - y + z = 5 ----- multiply by -1 -2x - y - z = -6 -------------- -x + y - z = -5 (result of multiplying by -1) -2x - y - z = -6 --------------add -3x - 2z = - 11 x - y + z = 5 ---- multiply by 3 3x + 3y - 2z = -5 ---------------- 3x - 3y + 3z = 15 (result of multiplying by 3) 3x + 3y - 2z = -5 ---------------add 6x + z = 10 -3x - 2z = -11 6x + z = 10 ---- multiply by 2 --------------- -3x - 2z = -11 12x + 2z = 20 (result of multiplying by 2) --------------add 9x = 9 x = 1 6x + z = 10 6(1) + z = 10 6 + z = 10 z = 10 - 6 z = 4 x - y + z = 5 1 - y + 4 = 5 5 - y = 5 -y = 0 y = 0 check... -2x - y - z = -6 -2(1) - 0 - 4 = -6 -2 - 4 = -6 -6 = -6 (correct) so x = 1, y = 0, and z = 4

OpenStudy (texaschic101):

It would probably be easier to use matrices....I am just not too good with those

OpenStudy (monkey*):

I was extremely confused, thank you so much for clearing it up for me. This is a huge help. Thank you so much! I'm not too great at matrices, either.